Suppose that it is determined that a $100 million portfolio could potentially lose $20 million once every 20 trading days What is the Var at 95% and 99% Confidence intervals?

Consider a portfolio of equities worth $1,000,000,000 with an expected daily return of 4% and a daily standard deviation of returns of 0.10%. Assume that all returns are normally distributed. What is the daily VaR on this portfolio at a tolerance threshold of 95%? What is the value at risk over a 27-day period?
1
Expert's answer
2020-06-16T12:36:53-0400

1."Var(95)=(1-0.95)\times20+1=2"

"Var(99)=(1-0.99)\times20+1=1.2"


2.

"VaR (95)=(-0.04+0.001\times1.645)\times1 000 000 000=-38 385 000"

"VaR(27)=−38385000\times\sqrt{27}=199454311"